OpenStudy (anonymous):

how do you solve this equation with a diode? circuit is in SERIES , made up of a 5v battery, its positive is connected to a 10ohm resistor,and connected to the negative side of the diode and then diode is connected to the positive side of an 8V battery and connected back to the NEGATIVE of the 5v battery? are they the same if in ideal model and ; how about at the 0.7 v model?

OpenStudy (08surya):

pls give circuit

OpenStudy (anonymous):

I guess you refer to the n-zone (cathode) and p-zone (anode) of the diode when you say negative side and positive side. When in a circuit, diodes have electrodes named anode and cathode. A description like the positive and negative sides of a diode is misleading as the sign of the voltage in its terminals is determined by the circuit itself. The circuit you describe can be solved assuming it is a Si diode and we can use the following model: \[I_D=\frac{ 8-5-V_D }{ 10 } \approx \frac{ 3-0.7 }{ 10 }=0.23A\]A more accurate calculation would be: \[10·I_D+V_D=3\rightarrow 10·I_S(e^{\frac{ V_D }{ V_T }}-1)+V_D=3\]which is a bit more complex to solve and requires that you know Is. This one can be solved with numerical methods or graphical methods